OpenStudy (anonymous):

The first number of three consecutive even integers equals the sum of the second and third. Find the three numbers. If x represents the smallest integer, then which of the following equations could be used to solve the problem? x = 2x + 6 x = 2x + 3 2x + 2 = x + 4

OpenStudy (anonymous):

these are my answer chooses x = 2x + 6 x = 2x + 3 2x + 2 = x + 4